Mr. H issues a 25 year mortgage of $225,000 at an annual interest rate of 4.5% to buy a house.The mortgage payments are made annually.

1.

What is Mr. H's annual payment of principal and interest?

$17,298

$13,353

$16,691

$15,174

2.

How much interest does Mr. H pay in the second year of the mortgage?

$11,283

$9,898

$8,710

$10,888

3.

Suppose that immediately after making the second annual payment, Mr. H has the opportunity to refinance the remaining mortgage balance at an annual rate of 3.5% for the remaining period of 23 years. What is the largest lump sum refinancing payment that he would be willing to make today to secure the lower cost financing? Assume that he continues to make annual payments on the new mortgage.

$22,345

$24,580

$19,664

$25,474

4.

Using the information from the refinancing question and assuming that Mr. H refinanced his mortgage at the lower rate after making two annual payments, how much is his remaining mortgage balance after making 3 of the lower annual payments in addition to the first two made before refinancing?

$171,886

$222,670

$195,325

$214,857
Answer #1

Formula sheet

A B C D E F G H I
2 1)
3 Calculation of annual Payment:
4
5 Annual payment can be calculated using PMT(RATE,NPER,PV,FV,TYPE) function in Excel as follows:
6
7 Given the following data:
8 Loan taken 225000
9 Duration of mortgage 25 Years
10 Interest rate 0.045
11
12 Annual Payment can be calculated as below:
13 RATE (Annual interst rate): =D10
14 NPER (No of Years): =D9
15 PV (Loan Amount): =-D8
16 FV 0
17 TYPE 0
18 Annual Payment =PMT(D13,D14,D15,D16,D17) =PMT(D13,D14,D15,D16,D17)
19
20 Hence Annual Payment is =D18
21 Thus fourth option is correct.
22
23 2)
24 Mortgage Table will be as follows:
25
26 Year Beginning Balance Annual Payment Interest Paid Principle Paid Ending Balance
27 1 =D8 =$D$20 =D27*$D$10 =E27-F27 =D27-G27
28 =C27+1 =H27 =$D$20 =D28*$D$10 =E28-F28 =D28-G28
29 =C28+1 =H28 =$D$20 =D29*$D$10 =E29-F29 =D29-G29
30 =C29+1 =H29 =$D$20 =D30*$D$10 =E30-F30 =D30-G30
31
32 Hence interest paid during second year is =F28
33 Thus second option is correct.
34
35 3)
36 The refinancing payment to be made on the existing mortgage will be the present value of mortgage.
37 Calculation of Value of Mortgage after2nd payment:
38
39 Value of mortage will be the present value of all the monthly payments.
40
41 Given the following data:
42 Annual Payments =D20
43 Interest Rate =D10
44 Number of years remaining =D9-2
45
46
47 Value of loan =Present value of annuity of monthly payments
48 =$15,173.78*(P/A,4.50%,23)
49 =D42*PV(D43,D44,-1,0) =D42*PV(D43,D44,-1,0)
50
51 Hence Value of loan after 96th payment is =D49
52
53 Calculation of Annual Payment after refinancing:
54
55 Monthly payment can be calculated using PMT(RATE,NPER,PV,FV,TYPE) function in Excel as follows:
56
57 Given the following data:
58 Loan taken =D51
59 Duration of mortgage =D44 Years
60 Interest rate 0.035
61
62 Annual Payment can be calculated as below:
63 RATE (Annual interst rate): =D60
64 NPER (No of Years): =D59
65 PV (Loan Amount): =-D58
66 FV 0
67 TYPE 0
68 Annual Payment =PMT(D63,D64,D65,D66,D67) =PMT(D63,D64,D65,D66,D67)
69
70 Hence Annual Payment is =D68
71
72 Calculation of profit or loss from Refinancing:
73
74 Savings on annual Payments =D20-D70 =D20-D70
75 Present Value of Savings =Annual Savings*(P/A,3.50%,23)
76 =D74*PV(D60,D59,-1,0) =D74*PV(D60,D59,-1,0)
77
78 Hence largest amount to be paid to secure refinancing is =D76
79 Thus first option is correct.
80
81 4)
82
83 Calculation of Value of refinanced Mortgage after 3 payment:
84
85 Value of mortage will be the present value of all the monthly payments.
86
87 Given the following data:
88 Annual Payments =D70
89 Interest Rate =D60
90 Number of years remaining =D59-3
91
92
93 Value of loan =Present value of annuity of monthly payments
94 =$13743.25*(P/A,3.50%,20)
95 =D88*PV(D89,D90,-1,0) =D88*PV(D89,D90,-1,0)
96
97 Hence Value of loan after 96th payment is =D95
98 Thus third option is correct.
99